# Query Performance Prediction using Relevance Judgments Generated by Large Language Models

### v1

@@ 2. Related Work
A majority of unsupervised and supervised post-retrieval QPP methods are proposed (Carmel and Yom-Tov, 2010) for predicting the performance of lexical rankers, such as query likelihood (Lafferty and Zhai, 2001), BM25 (Robertson et al., 2009). Unsupervised QPP methods can be classified into clarity-based (Cronen-Townsend et al., 2002), robustness-based (Zhou and Croft, 2007, 2006; Aslam and Pavlu, 2007), coherence-based (Arabzadeh et al., 2021a; Diaz, 2007), and score-based (Tao and Wu, 2014; Shtok et al., 2012; Cummins et al., 2011; Pérez-Iglesias and Araujo, 2010; Zhou and Croft, 2007). More recently, a set of supervised QPP methods have been proposed (Arabzadeh et al., 2021b; Hashemi et al., 2019; Zamani et al., 2018; Datta et al., 2022a, b; Chen et al., 2022; Khodabakhsh and Bagheri, 2023). NeuralQPP (Zamani et al., 2018) and Deep-QPP (Datta et al., 2022a) are optimized from scratch. NQA-QPP (Hashemi et al., 2019) and BERT-QPP (Arabzadeh et al., 2021b) fine-tune BERT (Devlin et al., 2019) to improve QPP effectiveness. Further, Datta et al. (2022c) propose qppBERT-PL, which considers list-wise-document information, while Chen et al. (2022) propose BERT-groupwise-QPP that considers both cross-query and -document information. Khodabakhsh and Bagheri (2023) propose a multi-task query performance prediction framework (M-QPPF), learning document ranking and QPP simultaneously.
However, studies have shown that post-retrieval QPP methods designed for lexical rankers struggle to predict the retrieval quality of neural rankers (Faggioli et al., 2023f; Hashemi et al., 2019), motivating several new unsupervised post-retrieval QPP methods designed for neural rankers. Datta et al. (2022b) propose a weighted relative information gain-based model (WRIG), which assesses a neural ranker for a given query by considering the relative difference of predicted performance between the given query and its variants; Zendel et al. (2023) assess a neural reranker by measuring the entropy of scores returned by it; Faggioli et al. (2023e) propose neural-ranker-specific ways of calculating regularization terms used by unsupervised post-retrieval QPP methods; vlachou2023coherence propose an unsupervised coherence-based QPP method that employs neural embedding representations to assess dense retrievers; and Singh et al. (2023) propose pairwise rank preference-based QPP (QPP-PRP) for predicting the performance of a neural ranker by measuring the degree to which a pairwise neural reranker (e.g., DuoT5 (Pradeep et al., 2021)) agrees with the ranked list returned by the neural ranker.
We present a novel QPP perspective: we start by automatically generating relevance judgments for the ranked list and then proceed to predict various IR evaluation measures. To the best of our knowledge, no prior work addresses QPP with a perspective similar to ours. Unlike regression-based QPP models (Khodabakhsh and Bagheri, 2023; Chen et al., 2022; Datta et al., 2022a; Arabzadeh et al., 2021b; Hashemi et al., 2019), which require training separate models to predict different IR evaluation measures, the training of LLMs for judging relevance in QPP-GenRE is agnostic to a specific metric, and different IR evaluation measures can be derived from the same set of generated relevance judgments. The closest work to QPP-GenRE, which is yet different, is qppBERT-PL (Datta et al., 2022c), in which they first predict the number of relevant items for each chunk in a ranked list and then aggregate those numbers into a general QPP score. However, the qppBERT-PL’s output is still presented as a single scalar, which is insufficient to accurately represent different evaluation measures; also, it is infeasible to predict any measures only using the number of relevant items in a ranked list.
@@
LLM for generating relevance judgments. Automatically generating relevance judgments is a long-standing goal in IR and has been studied over the past two decades (Makary et al., 2017, 2016; Ravana et al., 2015; Nuray and Can, 2006, 2003; Soboroff et al., 2001). Recent work shows promising results for using commercial LLMs (Thomas et al., 2023; Faggioli et al., 2023a; Gilardi et al., 2023) or pre-trained language models (MacAvaney and Soldaini, 2023) for automatically generating relevance judgments. Faggioli et al. (2023a) use zero-/few-shot prompting to instruct GPT-3.5 to predict the relevance of an item to a query. Similarly, Thomas et al. (2023) instruct GPT-4 by zero-shot prompting, and further add to the prompt a detailed query description and consider chain-of-thought (Wei et al., 2022). MacAvaney and Soldaini (2023) focus on a scenario where a relevant item for a given query is already known and use small-scale pre-trained language models (e.g., Flan-T5 (Chung et al., 2022) with 3B parameters) to estimate the relevance of another item to the query given the known relevant item. Some of the studies listed use LLM-generated relevance judgments for search system evaluation at a system level (Faggioli et al., 2023a; Thomas et al., 2023; MacAvaney and Soldaini, 2023) (i.e., the performance of ranked lists is aggregated to each system), and some at a query level (Thomas et al., 2023) (i.e., the performance of ranked lists is aggregated to each query). The evaluation is typically conducted offline as it requires different queries and multiple ranked lists for those queries (Faggioli et al., 2023a).
@@
Our work differs from the studies mentioned above: (i) unlike (Faggioli et al., 2023a; Thomas et al., 2023), we leverage open-source LLMs and further explore the impact of fine-tuning on the performance of open-source LLMs in generating relevance judgments; (ii) unlike (MacAvaney and Soldaini, 2023), we focus on a more practical scenario wherein no relevant item is known in advance for each query; and (iii) unlike (Faggioli et al., 2023a; Thomas et al., 2023; MacAvaney and Soldaini, 2023), which focus on applying generated relevance judgments to offline evaluation, we focus on QPP and predict the quality of a single ranked list for a query using LLM-generated relevance judgments, which the previous studies have not explored; QPP can be conducted either online (at query time) or offline (Faggioli et al., 2023d).
LLM for reranking. Recent studies on using LLMs for reranking have witnessed remarkable progress (Meng et al., 2024; Zhuang et al., 2023b; Bommasani et al., 2023; Ma et al., 2023a, b; Zhuang et al., 2023a; Drozdov et al., 2023; Sachan et al., 2022; Zhang et al., 2023b; Pradeep et al., 2023b; Tang et al., 2023; Pradeep et al., 2023a; Ma et al., 2023b; Sun et al., 2023b; Zhuang et al., 2023c; Hou et al., 2023). Those studies aim to predict the relevance order of items. The approaches closest to ours are some pointwise LLM-based rerankers that prompt an LLM to generate binary label tokens (e.g., “yes” or “no”) (Bommasani et al., 2023; Ma et al., 2023b) or fine-grained label tokens (Zhuang et al., 2023b) for a given query-item pair; they focus on using LLMs’ output logits of the label tokens as relevance scores to sort items in the descending of relevance. If we regard those label tokens as relevance labels, those rerankers (Bommasani et al., 2023; Ma et al., 2023b; Zhuang et al., 2023b) also have the potential to explicitly generate relevance judgments.
Our main contribution in this paper is the introduction of a novel QPP framework, which can be integrated with various approaches for judging relevance. We leave for future work the exploration of using those pointwise LLM-based rerankers as relevance judgment generators and integrating them with QPP-GenRE.
@@ 3. Task definition
This paper focuses on post-retrieval QPP (Carmel and Yom-Tov, 2010). Generally, a post-retrieval QPP method aims to estimate the retrieval quality of a ranked list with retrieved items induced by a ranker over a corpus in response to query without human-labeled relevance judgments, formally where indicates the predicted retrieval quality of the ranker in response to the query ; typically, is expected to be correlated with an IR evaluation measure, such as reciprocal rank (RR).
@@ 4.1. Overview of QPP-GenRE
We propose QPP-GenRE, which consists of two steps: (i) generating relevance judgments using LLMs, and (ii) predicting IR evaluation measures. In (i), we employ an LLM to generate relevance judgments for the top- retrieved items in the ranked list for a given query; to improve LLM’s effectiveness in generating relevance judgments, we train an LLM with PEFT using human-labeled relevance judgments. In (ii), we regard the generated relevance judgments as pseudo labels to calculate different IR evaluation measures.
@@ 4.2.1. Inference
where is a predicted relevance value for the item at rank . , where “1” indicates relevant and “0” irrelevant. We leave the prediction of multi-graded labels for the future. After automatically judging the top- items in the ranked list , we get a list of generated relevant judgments , where is the predicted relevance value for in . We design a prompt to instruct an LLM on the task of automatic generation of relevance judgments, illustrated in Figure 1.
@@
To further improve LLM’s effectiveness in generating relevance judgments, we use human-labeled relevance judgments to train an LLM with an effective PEFT method, QLoRA (Dettmers et al., 2023). Specifically, we first quantize an LLM model to 4-bit, add learnable low-rank adapters to each network layer of an LLM, and then optimize low-rank adapters. Formally, given the query and an item in the ranked list , we optimize the LLM to generate the human-labeled relevance value for the item . See Section 5 for more details.
@@ 4.3.1. Predicting precision-oriented measures
We compute a precision-oriented measure based on the LLM generated relevant judgments for the top- items in the ranked list . The following is an example to compute RR@k:
@@
4.3.2. An approximation strategy to predict recall-oriented measures
As the computation of a recall-oriented measure requires the information of all relevant items in the corpus for a given query , we need to automatically assess every item in corpus , which is infeasible due to the high computational cost. To address this issue, we devise an approximation strategy for predicting a recall-oriented IR measure, which only judges the top- items in the ranked list and uses the items predicted as relevant to approximate all relevant items in the corpus, to avoid the cost of judging the entire corpus. Fröbe et al. (2023); Moffat (2017); Lu et al. (2016) define normalized discounted cumulative gain (nDCG) (Järvelin and Kekäläinen, 2002) at a cutoff as a recall-oriented IR evaluation metric because it is normalized by a recall-oriented “best possible” ranking.111In this paper, we employ nDCG@10 and believe is a recall-oriented metric: Figure 2 illustrates that to reach saturation in predicting nDCG@10 values for ANCE and BM25, judgments up to the top 100 and 200 retrieved items are needed, respectively. If it were a precision-based metric, saturation could be achieved by judging around 10 items. nDCG@10 is also the most primary official IR evaluation metric in TREC-DL 19–22 (Craswell et al., 2022, 2021, 2020, 2019). Thus, we show an example of predicting nDCG@k (Järvelin and Kekäläinen, 2002), formally:
@@ 5. Experimental setup
Research questions. We address the following research questions:
@@
To what extent does QPP-GenRE improve QPP effectiveness for lexical and neural rankers in terms of a precision-oriented metric, RR@10, compared to state-of-the-art baselines?
@@
To what extent does QPP-GenRE improve QPP effectiveness for lexical and neural rankers in terms of a recall-oriented metric, nDCG@10, compared to state-of-the-art baselines?
@@
How deep do we need to automatically judge in a ranked list to effectively predict a recall-oriented IR evaluation measure?
@@
To what extent does fine-tuning LLaMA impact the quality of the generated relevance judgments and QPP effectiveness?
Datasets. We experiment with 4 widely-used IR datasets, TREC 2019–2022 deep learning (TREC-DL) tracks (Craswell et al., 2022, 2021, 2020, 2019). These datasets provide relevance judgments in multi-graded relevance scales per query. TREC-DL 19, 20, 21 and 22 have 43, 54, 53 and 76 queries, respectively. TREC-DL 19/20 and TREC-DL 21/22 are based on MS MARCO V1 (8.8 million passages) and V2 (over 138 million passages) passage ranking collections respectively.
Retrieval approaches. We consider BM25 (Robertson et al., 2009) as a lexical ranker, and ANCE (Xiong et al., 2021) as a neural-based dense retriever. To increase the comparability and reproducibility of our paper, we get the retrieval results of both rankers using the publicly available resource from Pyserini. We get BM25’s retrieval result with top-1000 retrieved items per query on TREC-DL 19–22 using the default parameters (, ); BM25’s actual nDCG@10 values are 0.506, 0.480, 0.446 and 0.269 on TREC-DL 19, 20, 21 and 22, respectively. We get ANCE’s retrieval result with top-1000 retrieved items per query on TREC-DL 19–20, using the publicly available dense vector index of ANCE on MS MARCO V1; ANCE’s actual nDCG@10 values are 0.645 and 0.646 on TREC-DL 19 and 20, respectively. We rely on the publicly available dense vector index of ANCE; at the time of writing, there is no dense vector index of ANCE publicly available on MS MARCO V2 for TREC-DL 21 and 22.333Building the dense vector index on MS MARCO V2 with over 138 million passages is resource-intensive and out of the scope of our work.
QPP baselines. We consider three groups of baselines: unsupervised post-retrieval QPP methods, supervised post-retrieval QPP methods, and the LLM-based QPP methods. The unsupervised QPP approaches that we consider are:
Weighted information gain (WIG) (Zhou and Croft, 2007) calculates the difference between retrieval scores of top- items in a ranked list and the retrieval score of the entire corpus.
Normalized query commitment (NQC) (Shtok et al., 2012) calculates the standard deviation of retrieval scores of top- items in a ranked list, normalized by the retrieval score of the entire corpus.
(Pérez-Iglesias and Araujo, 2010) it computes the standard deviation of retrieval scores from the first item to each point in a ranked list and outputs the maximum standard deviation.

### v2

@@ 2.2. Zero/few-shot prompting and parameter-efficient fine-tuning (PEFT) for LLMs
Automatically generating relevance judgments is a long-standing goal in IR that has been studied for multiple decades (Makary et al., 2017, 2016; Ravana et al., 2015; Nuray and Can, 2006, 2003; Soboroff et al., 2001). Recent studies have demonstrated promising results of using LLMs for the automatic generation of relevance judgments (Thomas et al., 2023; Faggioli et al., 2023a). In this paper we focus on studies into generating relevance judgments with discrete classes (e.g., “Relevant” or “Irrelevant”), instead of generating continuous relevance labels in real numbers (Yan et al., 2024). We discuss related studies into LLM-based automatic generation of relevance judgments from two dimensions: (i) how LLMs are used to generate relevance judgments, and (ii) their applications.
Recent studies have explored prompting commercial LLMs (e.g., GPT-3.5 and GPT 4) or open-source LLMs in zero- or few-shot manners. Specifically, Faggioli et al. (2023a) use zero- and few-shot prompting to instruct GPT-3.5 to predict the relevance of an item to a query. Thomas et al. (2023) instruct GPT-4 by zero-shot prompting, and add to the prompt a detailed query description and consider chain-of-thought (Wei et al., 2022). Similarly, Rahmani et al. (2024) follow Thomas et al. (2023)’s prompt to instruct GPT-4. Ma et al. (2024) instruct GPT-3.5 to generate relevance judgments for a domain-specific scenario, i.e., legal case retrieval (Ma et al., 2021); they use prompts specifically designed for this scenario. More recently, Upadhyay et al. (2024b) prompt GPT-4o in a zero-shot manner. Besides using commercial LLMs, only few studies (Zhang et al., 2024; Salemi and Zamani, 2024; Upadhyay et al., 2024a; Khramtsova et al., 2024) explore prompting open-source LLMs to generate relevance judgments. E.g., Khramtsova et al. (2024), Upadhyay et al. (2024a) and Salemi and Zamani (2024) prompt Flan-T5 (Chung et al., 2022), Vicuña-7B (Zheng et al., 2024) and Mistral (Jiang et al., 2023), respectively, in either zero-shot or few-shot manners. MacAvaney and Soldaini (2023) focus on a special scenario where a relevant item for a given query is already known and use Flan-T5 (Chung et al., 2022) to estimate the relevance of another item to the query given the known relevant item.
Recent studies have explored using LLM-generated relevance judgments to benefit (i) search system evaluation (Upadhyay et al., 2024a; Rahmani et al., 2024; Thomas et al., 2023; MacAvaney and Soldaini, 2023; Faggioli et al., 2023a), (ii) ranker selection (Khramtsova et al., 2024), (iii) item selection and retrieval quality evaluation in retrieval-augmented generation (RAG) (Zhang et al., 2024; Salemi and Zamani, 2024) and (iv) retriever fine-tuning (Ma et al., 2024). Concerning (i), recent studies (Rahmani et al., 2024; Upadhyay et al., 2024a; Abbasiantaeb et al., 2024; Faggioli et al., 2023a; Thomas et al., 2023; MacAvaney and Soldaini, 2023) explore evaluating search systems either entirely using LLM-generated relevance judgments or partially using LLM-generated relevance judgments (a.k.a. filling holes). They have demonstrated a high correlation between search system rankings based on LLM- and human-labelled relevance judgments. As to (ii), given a pool of dense retrievers, Khramtsova et al. (2024) select a suitable one for a target corpus by estimating their performance using LLM-generated queries and relevance judgments specific to the target corpus. For (iii), for item selection, Zhang et al. (2024) prompt LLMs to generate relevance judgments for retrieved candidate items in RAG; the items that are predicted as “relevant” are used for text generation. Zhang et al. (2024) observe that items selected via relevance prediction resulted in sub-optimal text generation quality. For retrieval quality evaluation, Salemi and Zamani (2024) generate relevance judgments for retrieved candidate items and aggregate those judgments into a score. However, Salemi and Zamani (2024) found that the aggregated score based on the LLM-generated relevance judgments achieves a low correlation with the text generation quality of RAG. Concerning (iv), Ma et al. (2024) fine-tune a legal case retriever on a training set augmented with LLM-generated relevance judgments. They show that fine-tuning a legal case retriever using the generated relevance judgments results in enhanced performance.
@@
Our work differs from the studies mentioned above: (i) we explore the possibility of fine-tuning open-source LLMs for generating relevance judgments; unlike MacAvaney and Soldaini (2023), we focus on a more practical scenario wherein no relevant item is known in advance for each query; and (ii) we focus on QPP and predict the ranking quality of a ranked list for a query using LLM-generated relevance judgments, which previous studies have not explored.
[…]
Recent studies on using LLMs for re-ranking have witnessed remarkable progress (Meng et al., 2024; Askari et al., 2023; Zhuang et al., 2023b; Bommasani et al., 2023; Ma et al., 2023a, b; Zhuang et al., 2023a; Drozdov et al., 2023; Sachan et al., 2022; Zhang et al., 2023c; Pradeep et al., 2023b; Tang et al., 2023; Pradeep et al., 2023a; Ma et al., 2023b; Sun et al., 2023b; Zhuang et al., 2023c; Hou et al., 2023). There are four paradigms of LLM-based re-ranking: pointwise, pairwise, listwise, and setwise (Zhuang et al., 2023c). Given a query, pointwise re-rankers produce a relevance score for each item independently, and the final ranking is formed by sorting items by relevance score (Ma et al., 2023a; Drozdov et al., 2023; Zhuang et al., 2023b; Sachan et al., 2022). The pairwise paradigm (Qin et al., 2023) eliminates the need for computing relevance scores; given a query and a pair of items, a pairwise re-ranker estimates whether one item is more relevant than the other for the query. Listwise re-rankers (Zhang et al., 2023c; Pradeep et al., 2023b; Tang et al., 2023; Pradeep et al., 2023a; Ma et al., 2023b; Sun et al., 2023b) frame re-ranking as a pure generation task and directly output the reordered ranked list given a query and a ranked list return by first-stage retriever (Zhang et al., 2023c; Pradeep et al., 2023b; Tang et al., 2023; Pradeep et al., 2023a; Ma et al., 2023b; Sun et al., 2023b). Given the low efficiency of pairwise (multiple inference passes) and listwise (multiple decoding steps) re-rankers, the setwise paradigm (Zhuang et al., 2023c) is meant to improve the efficiency while retaining re-ranking effectiveness. Given a query and set of items, an LLM is asked which item is the most relevant one to the query; these items are reordered according to the LLM’s output logits of each item being chosen as the most relevant item to the query, which only requires one decoding step of an LLM.
Our work differs from this line of research because we generate explicit relevance judgments with discrete classes (e.g., “Relevant” or “Irrelevant”), whereas studies into LLMs for re-ranking aim to predict the relevance order of items. However, using LLMs for generating relevance judgments and for re-ranking are intrinsically the same task: relevance prediction. Thus, an LLM-based re-ranker has the potential to serve as a relevance judgment generator.
Our main contribution in this paper is the introduction of QPP-GenRE, a novel QPP framework, which, in theory, can be integrated with various relevance prediction approaches. To demonstrate the compatibility of QPP-GenRE with various relevance prediction approaches, we adapt a state-of-the-art pointwise LLM-based re-ranker, RankLLaMA (Ma et al., 2023a), into a relevance judgment generator by applying a threshold for its re-ranking scores; we then integrate QPP-GenRE with this adapted RankLLaMA. Note that adapting other types of LLM-based re-rankers, e.g., pairwise and listwise LLM-based re-rankers into relevance judgment generators is challenging because they do not explicitly produce explicit relevance scores. Exploring their use as relevance judgment generators is beyond the scope of this paper.
@@ 3. Task definition
In this paper, we focus on post-retrieval QPP (Carmel and Yom-Tov, 2010). Generally, a post-retrieval QPP method aims to estimate the retrieval quality of a ranked list with retrieved items induced by a ranker over a corpus in response to query without human-labeled relevance judgments, formally:
where indicates the predicted retrieval quality of the ranker in response to the query ; typically, is expected to be correlated with an IR evaluation measure, such as reciprocal rank (RR).
@@ 4.1. Overview of QPP-GenRE
We propose QPP-GenRE, which consists of two steps: (i) generating relevance judgments using LLMs, and (ii) predicting IR evaluation measures. In (i), we employ an LLM to generate relevance judgments for the top- retrieved items in the ranked list for a given query; to improve LLM’s effectiveness in generating relevance judgments, we fine-tune an LLM with PEFT using human-labeled relevance judgments. In (ii), we regard the generated relevance judgments as pseudo labels to calculate different IR evaluation measures.
@@ 4.2.1. Inference
where is a predicted relevance value for the item at rank . , where “1” indicates relevant and “0” irrelevant. We leave the prediction of multi-graded labels as future work. After automatically judging the top- items in the ranked list , we get a list of generated relevant judgments , where is the predicted relevance value for in . We design a prompt to instruct an LLM on the task of automatic generation of relevance judgments, as illustrated in Figure 1.
@@
To further improve an LLM’s effectiveness in generating relevance judgments, we use human-labeled relevance judgments to train an LLM with an effective PEFT method, QLoRA (Dettmers et al., 2023). Specifically, we first quantize an LLM model to 4-bit, add learnable low-rank adapters to each network layer of the LLM, and then optimize low-rank adapters. Formally, given the query and an item in the ranked list , we optimize the LLM to generate the human-labeled relevance value for the item . See Section 5 for more details.
@@ 4.3.1. Predicting precision-oriented measures
We compute a precision-oriented measure based on LLM generated relevant judgments for the top- items in the ranked list . The following is an example to compute RR@k:
@@
4.3.2. An approximation strategy to predict measures considering recall
As the computation of a measure considering recall requires the information of all relevant items in the corpus for a given query , we need to automatically assess every item in corpus , which is infeasible due to the high computational cost. To address this issue, we devise an approximation strategy for predicting an IR measure considering recall, which only judges the top- items in the ranked list and uses the items predicted as relevant to approximate all relevant items in the corpus, to avoid the cost of judging the entire corpus. Fröbe et al. (2023); Moffat (2017); Lu et al. (2016) define normalized discounted cumulative gain (nDCG) (Järvelin and Kekäläinen, 2002) at a cutoff as a recall-oriented IR evaluation metric because it is normalized by a recall-oriented “best possible” ranking.111In this paper, we employ nDCG@10 and believe that nDCG@10 is a metric considering recall: Figure 3 illustrates that to reach saturation in predicting nDCG@10 values for ANCE and BM25, judgments up to the top 100 and 200 retrieved items are needed, respectively. If it were a precision-based metric, saturation could be achieved by judging around 10 items. nDCG@10 is also the most primary official IR evaluation metric in TREC-DL 19–22 (Craswell et al., 2022, 2021, 2020, 2019). Thus, here we show an example of predicting nDCG@ (Järvelin and Kekäläinen, 2002), formally:
@@ 5. Experimental setup
5.1. Research questions
In this section, we study the following research questions:
@@ 5.1. Research questions
To what extent does QPP-GenRE improve QPP quality for lexical and neural rankers in terms of RR@10, a precision-oriented IR metric, compared to state-of-the-art baselines?
@@
To what extent does QPP-GenRE improve QPP quality for lexical and neural rankers in terms of nDCG@10, an IR metric considering recall, compared to state-of-the-art baselines?
@@
How does judging depth in a ranked list affect the prediction of nDCG@10, an IR metric that considers recall?
@@
To what extent do fine-tuning and the choice of LLMs affect the quality of generated relevance judgments and QPP?
5.2. Datasets
We experiment with 4 widely-used IR datasets from the TREC 2019–2022 deep learning (TREC-DL) tracks (Craswell et al., 2022, 2021, 2020, 2019). These datasets provide relevance judgments in multi-graded relevance scales per query. TREC-DL 19, 20, 21 and 22 have 43, 54, 53 and 76 queries, respectively. TREC-DL 19/20 and TREC-DL 21/22 are based on the MS MARCO V1 and MS MARCO V2 passage ranking collections respectively. In the V1 edition, the corpus comprises 8.8 million passages while the V2 edition has over 138 million passages.
5.3. Retrieval approaches
We consider BM25 (Robertson et al., 2009) as a lexical ranker, and ANCE (Xiong et al., 2021) as a neural-based dense retriever. To increase the comparability and reproducibility of our paper, we get the retrieval results of both rankers using the publicly available resource from Pyserini (Lin et al., 2021). We get BM25’s retrieval result with top-1000 retrieved items per query on the TREC-DL 19–22 datasets using the default parameters (, ). BM25’s actual nDCG@10 values are 0.506, 0.480, 0.446 and 0.269 on TREC-DL 19, 20, 21 and 22, respectively. We get ANCE’s retrieval result with top-1000 retrieved items per query on TREC-DL 19–20, using the publicly available dense vector index of ANCE on MS MARCO V1. ANCE’s actual nDCG@10 values are 0.645 and 0.646 on TREC-DL 19 and 20, respectively. We rely on the publicly available dense vector index of ANCE; at the time of writing, there is no dense vector index of ANCE publicly available on MS MARCO V2 for TREC-DL 21 and 22.333Building the dense vector index on MS MARCO V2 with over 138 million passages is resource-intensive and out of the scope of our work.
5.4. QPP baselines
We consider three groups of baselines: unsupervised post-retrieval QPP methods, supervised post-retrieval QPP methods, and the LLM-based QPP methods. Specifically, we consider the following unsupervised QPP approaches that already showed high correlation with actual retrieval performance in previous work:
Weighted information gain (WIG) (Zhou and Croft, 2007) calculates the difference between retrieval scores of the top- items in a ranked list and the retrieval score of the entire corpus.
Normalized query commitment (NQC) (Shtok et al., 2012) calculates the standard deviation of retrieval scores of the top- items in a ranked list to a query; the standard deviation is normalized by the retrieval score of the entire corpus to the query.
(Pérez-Iglesias and Araujo, 2010) computes the standard deviation of retrieval scores from the first item to each point in a ranked list and outputs the maximum standard deviation.